what is the system of simultaneous equations by using the substitution method?
x=10-y
y=x-8

Answer by sofiyac(983)

x=10-y
y=x-8
Plug in the first equation into the second one and you get
y=10-y-8 combine like terms and you get
y=2-y now add y to each side and you get
2y=2 divide both sides by 2 and you get
y=1 now you plug that into the first original equaiton
x=10-1 then you simplify and you get
x=9
So the solution is (9,1)